OpenStudy (anonymous):

write an arithmetic sequence that has 3 arithmetic means between -1/4 and -19/12

OpenStudy (raden):

do u have options

OpenStudy (anonymous):

what do you mean?

OpenStudy (raden):

multiple choices.... i think there are many possible answers

OpenStudy (anonymous):

noo, no multiple choice i have to write 3 numbers between -1/4 and -19/12 what is the sequence between them ?

OpenStudy (raden):

the 1st possible sequence is -18/12, -17/12, -16/12 the 2nd, -18/12, -16/12, -14/12 etc..

OpenStudy (anonymous):

umm i think it wants the numbers between =1/4 and -19/24

OpenStudy (raden):

why like that, give ur suggestion

OpenStudy (anonymous):

hold on i think i got it, thanks anyways :)

OpenStudy (raden):

yw, what's ur answer

OpenStudy (anonymous):

I have to use this formula an = a1 + (n – 1)d

OpenStudy (raden):

yeah, i know that... i have used above

OpenStudy (anonymous):

ok so far i have putted in this -19/12 = -1/4 + (n - 1)d i dk n and d

OpenStudy (anonymous):

i dont know what to put in n and d

OpenStudy (anonymous):

will it be -19/12 = -1/4 + (6 - 1)d ? so what is the difference? what do i do to find it?

OpenStudy (anonymous):

ok so far -19/12 = -1/4 + 5d is that correct?

OpenStudy (anonymous):

so i got d = -4/15

OpenStudy (anonymous):

??

OpenStudy (raden):

actually, "between" means that the number more than under interval and less than up interval... ex. between 2 and 10, so 2 < x < 10

OpenStudy (anonymous):

ok

OpenStudy (raden):

so ur numbers between -1/4 and -19/12 or between -3/12 and -19/12 it means -19/12 < the numbers < -3/12

OpenStudy (anonymous):

umm ok

OpenStudy (raden):

that's why i said many possible answer... :) sorry if my opini is wrong

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

hey so will this be right? 19/12 -(-1/4) = -19/12 + 1/4 = -19/12 + 3/12 = -16/12 = -4. Since the difference is -4, divide by 4, = -1.